# http://www.michaeltimothybennett.com/

### v2

@@ 3.1.2 Why confounding tends to occur:
We now briefly argue that abstraction is goal directed, meaning the tasks an abstraction layer tends to represent are those it is best suited to represent, and that this implies weak constraints take simple forms. There are several reasons to believe the formation of abstraction layer is goal directed, depending upon the context in which we consider complexity. In the case of a computer, a human has specifically designed each abstraction layer to express that which is needed for a purpose. What separates x86 from a higher level abstraction layer like Numpy is that the former has a more general intended purpose, expressing “weaker” constraints. We tend to construct abstraction layers to be as versatile as possible whilst satisfying a particular need. More generally natural selection favours adaptation, which means generalisation, which is maximised by preferring weaker policies [17]. Biological cognition is not limited to the brain [26], meaning the mind is not neatly confined within a well defined neurological abstraction layer. Instead the multiscale competency architectures observed in living organisms [27] amount to self organising abstraction layers. Because natural selection favours adaptable organisms, these abstraction layers will be selected to represent the weakest policies which constitute fit behaviour (a weaker policy is more adaptable). More generally, we speculate that phase transitions motivate the emergence of self preserving goal directed behaviour, by destroying some physical structures and preserving others. Such goal directed abstraction must minimise the size of vocabularies at higher levels, whilst also maximising the weakness of the policies they can express (two opposing pressures). This is because a larger a vocabulary exponentially increases the space of outputs and policies [28], which may conflict with with finite time and space constraints. A larger vocabulary would make inference and learning less tractable (more “complex” in the sense of being a more difficult search problem that takes up more time). To maximise the weakness of policies in higher levels of abstraction, while minimising the size of the vocabulary in which they’re expressed, weaker policies must take simpler forms.

### v3

@@ 3.1.2 Why confounding tends to occur:
We now briefly argue that abstraction is goal directed. This means the tasks an abstraction layer tends to represent are those it is best suited to represent, which implies weak constraints take simple forms. There are several reasons an abstraction layer is biased toward particular goals, depending upon the context in which we consider complexity. In the case of a computer, a human has specifically designed each abstraction layer to express that which is needed for a purpose. What separates x86 from a higher level abstraction layer like Numpy is that the former has a more general intended purpose, expressing “weaker” constraints. We tend to construct abstraction layers to be as versatile as possible whilst satisfying a particular need. More generally natural selection favours adaptation, which means generalisation, which is maximised by preferring weaker policies [bennett2023b]. Biological cognition is not limited to the brain [ciaunica2023], meaning the mind is not neatly confined within a well defined neurological abstraction layer. Instead the multiscale competency architectures observed in living organisms [levin2024] amount to self organising abstraction layers. Because natural selection favours adaptable organisms, these abstraction layers will be selected to represent the weakest policies which constitute fit behaviour (a weaker policy is more adaptable). More generally, we speculate that phase transitions motivate the emergence of self preserving goal directed behaviour, by destroying some physical structures and preserving others. Such goal directed abstraction must minimise the size of vocabularies at higher levels, whilst also maximising the weakness of the policies they can express (two opposing pressures). This is because a larger a vocabulary exponentially increases the space of outputs and policies [bennett2022a], which may conflict with with finite time and space constraints. A larger vocabulary would make inference and learning less tractable (more “complex” in the sense of being a more difficult search problem that takes up more time). To maximise the weakness of policies in higher levels of abstraction, while minimising the size of the vocabulary in which they’re expressed, weaker policies must take simpler forms.
[…]